Find the value of a for which the polynomial x^5-ax²-ax+1has −1 as a root with multiplicity at least 2.
1
Expert's answer
2020-03-06T11:39:40-0500

it is possible if a = -5. Indeed, x5-ax2-ax+1=(x+1)(x4-x3+x2-x+1)-ax(x+1)=(x+1)(x4-x3+x2-x+1-ax)=(x+1)(x+1)(x3-2x2+3x+1)